# Truthful Text Sanitization Guided by Inference Attacks

### v1

@@ 6 Conclusions and future work
We presented INTACT, a two-stage approach to automatically sanitize disclosive text spans in documents by leveraging instruction-tuned large language models. In contrast with related works, our approach is not limited by the narrow scope of structured knowledge bases and focuses on generating truthful replacements, which are of outmost importance in secondary uses of the data, such as medical research, insurance or law. We ensure the replicability of our method by using open-weights LLMs, which can also be run locally to guarantee private and secure handling of sensitive data. Empirical results show that our method provides the best privacy/utility trade-off, while avoiding distorting the original semantic content of the document as other LLM-based approaches do.
We have also proposed automatic utility and privacy metrics for text anonymization, which capture the notion of semantics and disclosure risk better than standard precision and recall metrics, and do not require (costly) manual annotations as ground truth.
In future work, we aim to enhance both the ordering of replacement candidates and their grammatical integration within sentences. Even though a grammatically correct output is specifically requested in the replacement generation prompt, some replacements would still produce ungrammatical phrases (e.g., double prepositions like in in) if inserted, partly due to inconsistencies in the inclusion of these within the manually annotated spans of the dataset used. Additionally, while the replacement suggestions generally range from the most specific to the most abstract, the ordering isn’t always consistent. We also plan to explore additional methods for matching the original spans against those guessed by the LLM. These enhancements could potentially be implemented through additional LLM steps to further improve the quality of the sanitized text.
@@ Acknowledgments
We acknowledge support from the Norwegian Research Council (CLEANUP project (http://cleanup.nr.no/), grant nr. 308904), the Government of Catalonia (ICREA Acadèmia Prize to D. Sánchez, and grant 2021SGR-00115), MCIN/AEI/ 10.13039/501100011033 and “ERDF A way of making Europe” under grant PID2021-123637NB-I00 “CURLING”, and the EU’s NextGenerationEU/PRTR via INCIBE (project “HERMES” and INCIBE-URV cybersecurity chair).
@@
During the preparation of this work the author(s) used ChatGPT (with GPT-4o) in order to improve readability and fluency in parts of the text. After using this tool/service, the author(s) reviewed and edited the content as needed and take(s) full responsibility for the content of the published article.
@@ A.1 Replacement Generation
The prompt used for generating replacement candidates in a one-shot setting is presented in Table 3. It is set up as a multi-turn chat to accommodate the example (listed Table 2) of a matching entity type compared to the target span.
Table 3: One-shot prompt for replacement generation.
Role Message
Here is an example:
{Example sentence context}
Sorted replacements for {[[Example target span]]}:
Assistant {Example of sorted replacements}
User Now do the same for the sentence below.
Original:
{Sentence context}
Sorted replacements for {[[Target span]]}:
A.2 Replacement Selection
Table 4 shows the one-shot prompt used for the guessed-based selection of replacement candidates. Here, the same example was used in all cases.
Table 4: Prompt for guessing a PII span based on a replacement candidate and the rest of the text with replacements instead of PIIs.
Role Message
Text: “early 1980s” $\rangle$ Guess: “3 September 1981”. List more likely guesses first. Do not return anything else. See the example below.
Guesses for “a European Research Institute”:
- the Institute of Polish Literature, University of Warsaw
- the Center for European Neighborhood Studies
- the Centre for Research on the History and Culture of Eastern Europe
- the Leibniz Institute for the History and Culture of Eastern Europe
User Do the same for the text below.
Text: {Text with replacements}
Guesses for {[[Target span]]}:
Table 5 shows statistics over entity type replacement.

### v2

@@ 5.7 Truthfulness and abstraction level
5.8 Generalizability to other LLMs
To empirically validate the generalizability of our method to other LLMs, we applied it to the TAB test set using another open-weights model, namely Llama 3.1 8B Instruct [^17] https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct [25]. Although both Llama 3.1 and Mistral 7B are transformer-based, they have some architectural differences, with Mistral 7B emphasizing efficiency through optimized attention mechanisms such as Sliding Window Attention. However, while Mistral 7B is primarily designed for English, Llama 3.1 offers strong performance in seven languages.
For the Llama 3.1 experiments, the prompts, parameter settings, and infrastructure were identical to those used in the Mistral-based experiments described in Section 5.2. We evaluated the outcomes using the same automatic utility and privacy metrics as in Sections 5.5 and 5.6, with one change: the background knowledge used by TRIR was adapted to this specific data subset. In Section 5.6, the background knowledge was a superset twice the size of the protected data. To replicate an equivalent level of adversarial robustness, we used the background knowledge from TAB’s test and dev sets. Since both sets have the same size (i.e., 127 individuals/documents), this forms a superset twice the size of the protected data, maintaining a comparable threat model. Table 3 reports the results for both Mistral and Llama LLMs, along with the suppression-based baseline and related work for reference on the TAB test set.
Table 3: Utility and privacy evaluation for LLM generalizability on the TAB test subset with Llama 3.1 8B. ($\uparrow$ for higher is better, $\downarrow$ for lower is better)
[…]
Suppression 77.7% 30.2% 11.0%
Entity type 84.8% 64.2% 13.4%
Presidio [39] 85.8% 66.0% 11.8%
Dou et al. [18] 88.4% 58.7% 15.0%
Mistral INTACT 88.5% 80.0% 12.6%
Mistral least-specific 88.5% 57.7% 13.4%
Mistral most-specific 91.3% 89.9% 14.2%
Llama INTACT 88.7% 86.6% 13.4%
Llama least-specific 87.4% 56.6% 13.4%
Llama most-specific 90.4% 88.7% 15.0%
Results demonstrate that INTACT performs consistently across both LLMs, confirming its generalizability. Llama 3.1 achieves slightly better sanitization performance than Mistral 7B, with a NMI 6% higher while maintaining TPS and obtaining a only 1% higher TRIR –which, for this 127 individuals set, implies that only one additional individual has been correctly re-identified. In addition, the superior privacy and utility trade-off observed for INTACT in previous sections also holds here: INTACT always achieves the highest NMI, equal or higher TPS, and equal or lower TRIR compared to related work, exceeding suppression-based masking by only 3%. For least-specific and most-specific replacements, Llama’s scores are about one percentage point below Mistral’s. This suggests that Llama produced slightly inferior lists of sorted replacement candidates in the generation step, but the gap is small compared to the differences with other methods. Furthermore, it highlights the impact of the Llama-based selection step, which, despite starting from worse candidate lists, still produced outcomes on par with –or better in the case of NMI– than Mistral-based INTACT. Mistral and Llama differ in how often the entity type was selected instead of the model-generated replacements. For Llama, the entity type was chosen ca. 5% less frequently, with a corresponding increase in the selection of the second most-specific candidate (selected for 12.8% of entity mentions for Mistral vs. 17.8% for Llama, see Figures 8 and 9 in A).
6 Discussion
The evaluation results from the previous sections corroborate that the identified gap in the literature – namely, the absence of a principled method for generating truthful and utility-preserving replacements for sensitive text spans – is effectively addressed by the proposed INTACT approach.
Our utility analysis shows that while INTACT performs on par with related work in intrinsic evaluation (TPS), it substantially outperforms it in extrinsic evaluation, achieving over a 10% improvement in NMI on TAB. This extrinsic metric is particularly relevant for practitioners, as it reflects performance in a generic downstream task. Accordingly, we can claim that INTACT achieves effective utility preservation.
This utility gain comes without a notable reduction in privacy protection. Privacy was assessed using the TRIR metric under a realistic threat model –one that specifies plausible background knowledge and computational resources (see Section 5.6). In this way, our adversarial robustness test aligns with Recital 26 of the GDPR [^18] https://gdpr-info.eu/recitals/no-26/ [24], which states that anonymous data falls outside the scope of the regulation and that ensuring anonymity requires considering the resources likely to be used for re-identification. The empirical re-identification risk for INTACT was not substantially higher than that of redaction (i.e., suppression-based masking). Concretely, the risk figure was approximately 9% (equivalent to $k=11$ or greater in probabilistic $k$-anonymity terms [50, 59]), what according to [20] corresponds to a strong degree of anonymity, as attested by two independent health regulators [21, 28].
Our manual evaluation of truthfulness and abstraction further supports these findings: INTACT’s replacements preserved the core meaning of original spans in 93% of cases, while avoiding simple rewrites in 85% of cases. This indicates that the method produces genuine generalizations, which can improve utility without significantly undermining privacy.
Importantly, these results were achieved using relatively small, instruction-tuned LLMs with open weights—specifically, Mistral (7B parameters) and Llama 3.1 (8B parameters). This demonstrates that compact models can be highly effective for efficient text sanitization. The strong performance of Llama 3.1 further confirms that the method generalizes well across different model architectures, with no notable loss in either utility preservation or privacy protection. This consistency suggests that INTACT’s effectiveness is rooted in its core design principles rather than in model-specific characteristics, highlighting its robustness and broad applicability.
Overall, INTACT presents a compelling solution for data holders aiming to release de-identified documents in compliance with privacy regulations, while maintaining the semantic and analytical integrity of the data. However, integrating the proposed method into end-to-end NLP systems must take the three following aspects in consideration:
  • 1.
    To support languages beyond English, the selected LLM [^19] This also holds for the generalization heuristics employed for DATETIME entities. must support the target language to accurately identify and replace sensitive spans.
  • 2.
    Due to the computational costs of running inference attacks on candidate replacements, the current approach is best suited for small- to medium-size document collections. Future improvements in data handling and candidate selection strategies could significantly enhance runtime performance and reduce computational overhead.
  • 3.
    Finally, like all sanitization techniques, INTACT does not offer ex ante privacy guarantees. Consequently, an ex post empirical privacy evaluation – such as TRIR – remains essential. Complementary assessments of utility preservation (e.g., TPS or NMI), truthfulness, and abstraction are also important to ensure the quality of the released documents. When these evaluations themselves rely on language models, it is important to ensure those support the target language; for instance, multilingual sentence-transformers with aligned cross-lingual representations can be used for TPS in multilingual contexts.
@@ 7 Conclusions and future work
We presented INTACT, a two-stage approach to automatically sanitize disclosive text spans in documents by leveraging instruction-tuned large language models. In contrast to previous approaches, the presented method is not limited by the narrow scope of structured knowledge bases and focuses on generating truthful replacements, which are of outmost importance in secondary uses of the data, such as medical research, insurance or law. A key benefit of the presented method is its ability to identify the most privacy-preserving candidates by testing whether a language model can infer the original span through an inference attack. Given the advanced capabilities of LLMs as adversaries (see e.g. [12]), this helps ensure that replacements duly conceal the personally identifiable information present in the original text spans. We ensure the replicability of our method by using open-weights LLMs, which can also be run locally to guarantee private and secure handling of sensitive data. Empirical results show that our method provides a better privacy/utility trade-off and avoids distorting the original semantic content of the document as other LLM-based approaches do.
We have also proposed Text Preserved Similarity (TPS), an automatic metric for text anonymization that captures the notion of semantics better than standard precision metrics, does not require (costly) manual annotations as ground truth and is compatible with sanitization.
In future work, we aim to enhance both the ordering of replacement candidates and their grammatical integration within sentences. Even though a grammatically correct output is specifically requested in the replacement generation prompt, some replacements occasionally produce ungrammatical phrases (e.g., double prepositions like in in) if inserted, partly due to inconsistencies in the inclusion of these within the manually annotated spans of the dataset used. The fluency of the output is an additional aspect that could be assessed manually by future work. Moreover, while the replacement suggestions generally range from the most specific to the most abstract, the ordering is not always consistent. We also plan to explore additional methods for matching the original spans against those guessed by the LLM. These enhancements could potentially be implemented through additional LLM steps to further improve the quality of the sanitized text. Future studies could also investigate whether these challenges persist across other LLMs. Given that some of our findings may be limited to the specific LLMs tested, a broader comparative analysis based on a variety of model types, sizes and configurations would allow for validating the robustness and generalizability of the proposed approach and results.
Future work will also seek to address the limitation of TPS with assessing terms’ similarity in isolation, mentioned in Section 5.5. Tackling this limitation requires a shift from term-level to context-aware similarity metrics, such as those based on sentence-level or passage-level embeddings. While such embeddings offer a more holistic representation by capturing the semantics of entire sentences, their integration into the TPS framework poses methodological challenges. Specifically, the similarity computation would no longer be a function solely of the original and replacement spans, but would instead need to consider the full textual environment, including potentially subtle changes in meaning or tone. This shift thus requires a revised scoring that can meaningfully interpret and weigh contextual information.
A final limitation of our work is the use of a benchmark limited to English texts and focusing on the legal domain, which may constrain generalizability across domains and languages. We will therefore seek to extend the evaluation to other domains and explore the use of multilingual LLMs to assess the method’s broader applicability.
@@ Acknowledgments
We acknowledge support from the Norwegian Research Council (CLEANUP project (http://cleanup.nr.no/), grant nr. 308904), the Government of Catalonia (ICREA Acadèmia Prize to D. Sánchez, and grant 2021SGR-00115), MCIN/AEI/ 10.13039/501100011033 and “ERDF A way of making Europe” under grant PID2021-123637NB-I00 “CURLING”, and the EU’s NextGenerationEU/PRTR via INCIBE (project “HERMES” and INCIBE-URV cybersecurity chair). B. Manzanares-Salor is also supported by the Spanish Government under an FPU grant (ref. FPU23/01785).
@@
During the preparation of this work the author(s) used ChatGPT and Microsoft Co-pilot (with GPT-4o) in order to improve readability and fluency in parts of the text. After using this tool/service, the author(s) reviewed and edited the content as needed and take(s) full responsibility for the content of the published article.
@@ Appendix A LLM-based text sanitization
Table 5 contains statistics over entity type replacement with the Mistral model.
@@
Figures 8 and 9 show how often replacement candidates with different levels of specificity were chosen by the Mistral and the Llama models respectively for the TAB test set. When all five candidates were predicted by the inference attacks, Entity Type is chosen (e.g.PERSON).
Figure 8: Frequency of INTACT’s replacement choice for the test set of TAB with the Mistral 7B Instruct model.
Figure 9: Frequency of INTACT’s replacement choice for the test set of TAB with the Llama 3.1 8B Instruct model.